# Joint Source-Channel Optimization for UAV Video Coding and Transmission

### v2

@@ Abstract
This paper is concerned with unmanned aerial vehicle (UAV) video coding and transmission in scenarios such as emergency rescue and environmental monitoring. Unlike existing methods of modeling UAV video source coding and channel transmission separately, we investigate the joint source-channel optimization issue for video coding and transmission. Particularly, we design eight-dimensional delay-power-rate-distortion models in terms of source coding and channel transmission and characterize the correlation between video coding and transmission, with which a joint source-channel optimization problem is formulated. Its objective is to minimize end-to-end distortion and UAV power consumption by optimizing fine-grained parameters related to UAV video coding and transmission. This problem is confirmed to be a challenging sequential-decision and non-convex optimization problem. We therefore decompose it into a family of repeated optimization problems by Lyapunov optimization and design an approximate convex optimization scheme with provable performance guarantees to tackle these problems. Based on the theoretical transformation, we propose a Lyapunov repeated iteration (LyaRI) algorithm. Both objective and subjective experiments are conducted to comprehensively evaluate the performance of LyaRI. The results indicate that, compared to its counterparts, LyaRI achieves better video quality and stability performance, with a 47.74% reduction in the variance of the obtained encoding bitrate.
@@ I Introduction
Benefiting from their distinctive advantages of low cost and high mobility unmanned aerial vehicles (UAVs) have been increasingly adopted in more and more practical scenarios, such as emergency search and rescue, geological disaster monitoring, and forest fire detection. Compared with platforms including satellites and near-space airships, UAVs can quickly respond to emergency relief and earth observation missions, achieving focused coverage of target areas [1]. Additionally, as the terminal nodes of the integrated space-air-ground information network, UAVs also undertake functions such as information perception and transmission, which can enhance local information service capabilities for target areas [2, 3].
In the application of UAVs, video acquisition and transmission constitute one of their pivotal missions. UAVs are capable of capturing targets through on-board cameras and subsequently transmitting coded and compressed videos via on-board data transmission modules. The prompt acquisition of on-site dynamic information via UAV videos can secure invaluable time for emergency rescue operations. As a result, the UAV video coding and transmission technology has emerged as a focal point of interest within both academia and industry [4, 5, 6].
However, the timely and efficient transmission of UAV video streams is fraught with considerable challenges [7]. Firstly, a UAV network is bandwidth-constrained while the video is characterized by its substantial information and stringent timeliness. It is essential to mitigate information redundancy through video coding for UAV video transmission. Nevertheless, the computational complexity and energy expenditure associated with video coding are exceedingly high, presenting a significant challenge for UAVs that are subject to severe energy constraints. Secondly, the occurrence of congestion and packet loss, potentially leading to visual artifacts or stalling in video decoding at the receiver, may arise if the video coding bitrate surpasses the UAV channel capacity [8]. Moreover, influenced by factors such as mobile positioning, signal interference, and rapid channel fading, UAV channel capacity is subject to dynamic fluctuations. Consequently, ensuring the match of UAV video coding bitrate with UAV channel capacity is a formidable challenge.
@@ I-A State of the Arts
All these studies assumed unlimited encoding time and power consumption. They also focused on the selection of video coding quantization parameters based on the R-D relationship and aimed to achieve the best encoding performance. However, the time and power consumption of video coding have been confirmed to significantly impact the video bitrate and distortion. In [15], the authors explored semi-extreme sparse coding of transformed residuals at medium and low bitrates to enhance video compression efficiency. The work in [16] indicated that video bitrate and distortion were primarily determined by the distribution characteristics of the transformed residuals. Transformed residuals were mainly influenced by motion estimation (ME) and quantization parameters. The accuracy of ME was closely related to the sum of absolute differences (SAD) of macroblocks. The number of calculations of macroblock SAD was proportional to the encoding complexity. Considering that the time and power consumption of video coding were increasing functions of encoding complexity, the authors in [16] deduced a quantitative relationship between video coding delay, power, rate, and distortion, and extended the traditional R-D model to a delay-power-rate-distortion (d-P-R-D) model. Based on this model, a new rate control (RC) method was proposed, aiming to minimize encoding distortion under constraints of rate, delay, and power. However, there are two aspects for improvement in this research. Firstly, the work is conducted for the earlier video coding standard advanced video coding (AVC). High efficiency video coding (HEVC) and versatile video coding (VVC), as the subsequent video coding standards[17], offer lower coding bitrates under the same video quality conditions. Therefore, it is necessary to further investigate the d-P-R-D model for the HEVC standard. Secondly, although the end-to-end delay composition of video transmission systems was analyzed in [16], the proposed d-P-R-D model was only for video source coding. To achieve better end-to-end video transmission QoE, it is essential to design a joint source-channel video coding and transmission scheme.
@@
From the perspective of joint source-channel design, efficient video source coding can reduce the video bitrate while maintaining the same video quality. The reduction in video bitrate leads to decreasing transmit power consumption and transmission delay. However, efficient video source coding implies higher computational complexity, which result in increased power consumption and delay in video coding. Given the constraints of end-to-end latency, increasing encoding latency to achieve better compression performance may reduce the transmission latency budget, subsequently increasing the risk of distortion during transmission. Similarly, changes in encoding power consumption can affect the power consumption of video transmission under a given power constraint of video transmission system. Researchers have extensively studied joint source-channel optimization [18, 19, 20, 21, 22, 23, 24, 25]. For instance, the work in [18] proposed a new scheme for dynamically adjusting the bitrate based on channel signal-to-noise ratio (SNR) and image content. In [19], the authors presented a joint source-channel adaptive RC scheme for wireless image transmission based on image feature mapping, entropy awareness, and channel conditions. The authors in [21] introduced a novel video codec structure based on compressed sensing. They also investigated the minimization of system power consumption by jointly optimizing video coding bitrate and multi-path transmission bitrate under distortion and energy consumption constraints. The work in [23] investigated the trade-off between latency, power, and distortion and achieved deep cross-layer optimization of video coding, queuing control, and wireless transmission through a joint lossy compression and power allocation scheme. However, the aforementioned research focused on the field of wireless video transmission and did not specifically target UAV video transmission. To address this, the authors in [8] conducted cross-layer design research on UAV-based video streaming transmission. They proposed a quality of service (QoS) strategy that dynamically adjusted the transmission mode according to network load, latency, and packet loss rate. Nevertheless, they focused on the overall system design of multi-link parallel transmission, without delving into the issue of UAV video coding. Considering the dynamic nature and power constraints of UAVs, it is necessary to conduct joint resource allocation and control for UAV video coding and transmission under the constraints of end-to-end latency and power consumption. In this way, the QoE of video transmission at the receiving end can be improved.
@@ II-A AtG Channel Model
This paper primarily considers the application scenarios where UAVs are deployed for emergency search and rescue, geological disaster monitoring, forest fire detection, and other missions that involve encoding and transmitting videos in a timely manner. Particularly, the scenario investigated in this paper encompasses a UAV and an emergency communication vehicle (ECV). Given the potential for road traffic disruptions in areas such as mountainous and post-disaster zones, the ECV is deployed in proximity to the affected region to launch the UAV. The UAV is tasked with conducting surveillance and reconnaissance missions, capturing video footage for transmission, as illustrated in Fig. 1. The captured video is encoded and compressed by the UAV and then promptly delivered to the ECV. To facilitate the construction of a mathematical model for the video transmission process, the temporal domain is discretized into a sequence of time slots, denoted by $t=\{1,2,\ldots\}$. Due to its limited surveillance coverage range, the UAV is deployed to fly along a circular trajectory to extend the coverage range. The two-dimensional (2-D) horizontal coordinate of the UAV at time slot $t$ is represented as $\bm{q_{uav}}\left(t\right)={[{x_{uav}}\left(t\right),{y_{uav}}\left(t\right)]^{\rm{T}}}$, and the position of the ECV is denoted as $\bm{q_{ecv}}={[{x_{ecv}},{y_{ecv}}]^{\rm{T}}}$. Consider that the UAV flies at a constant altitude $H$, and the altitude of the ECV is negligible compared to $H$. The horizontal projection distance between the UAV and the ECV at time slot $t$ can be expressed as $dis(t)=\sqrt{{{({x_{uav}}\left(t\right)-{x_{user}})}^{2}}+{{({y_{ecv}}\left(t\right)-{y_{ecv}})}^{2}}}$.
@@ II-B1 Video Bitrate and Distortion Models
The research presented in this paper is based on the HEVC standard. Although HEVC continues to employ the traditional block-based hybrid coding framework, which includes key modules such as prediction, transformation, quantization, entropy coding, and loop filtering, it introduces innovative coding techniques in each of these modules. These techniques include quadtree-based block partitioning, up to $35$ intra-prediction modes, inter-frame variable-sized discrete cosine transform (DCT), context-based adaptive binary arithmetic coding (CABAC), and sample adaptive offset (SAO) for loop filter compensation. Compared to AVC, HEVC significantly improves compression efficiency, achieving an approximately $50\%$ reduction in bitrate while maintaining the same video quality, making it highly suitable for video transmission in bandwidth-constrained networks. When compared to the next-generation video coding standard, e.g., versatile video coding (VVC), HEVC demonstrates a higher energy efficiency ratio [17]. Therefore, this paper constructs a d-P-R-D model for video coding based on HEVC and adopts the IPPPP coding mode without loss of generality.
@@
In the IPPPP coding mode, both encoding bitrate and distortion of P-frames in block-based hybrid video coding can be approximately represented as functions of the standard deviation of transformed residuals and the quantization step $Q$ [16]. Based on the research findings in literature[27], the functional expression $\sigma$ can be obtained through nonlinear regression analysis. This paper selects standard video test sequences CITY (CIF, 352*288) and Coastguard (CIF, 352*288) and employs the HM-16.20 encoder to obtain encoding data samples. In the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, since quantization parameter $QP$, search range $\lambda$, and reference frames $\chi$ are independent parameters, their individual impacts on $\sigma$ can be assessed separately. It should be noted that there is a one-to-one mapping relationship between the quantization parameter $QP$ mentioned in HM-16.20 and the quantization step $Q$ involved in the model. This paper first fixes $QP$ and $\chi$, and changes the value of $\lambda$ to obtain transformed residual data and calculate $\sigma$. Through nonlinear regression analysis, it has been found that $\sigma$ is approximately exponentially related to $\lambda$. Secondly, by fixing $\lambda$ and $\chi$, and changing $QP$, it can be concluded that $\sigma$ is approximately linearly related to $QP$. Additionally, it is observed that the rate of change of $\sigma$ versus $\chi$ is much smaller than that of $\sigma$ versus $QP$, as well as $\sigma$ versus $\lambda$. It indicates that $\chi$ has a relatively smaller impact on $\sigma$ compared to $QP$ and $\lambda$. To reduce the computational complexity, this paper assigns a given value ${\chi_{0}}=1$ to $\chi$[16]. Therefore, this paper adopts the following expression to calculate $\sigma$
@@ II-B2 Coding Delay and Power Models
In the research field of video coding, encoding complexity is an important indicator of quantization encoding delay. Compared to AVC, HEVC employs flexible coding units (CU) partitioning based on a quadtree structure and a richer set of coding modes. It results in relatively higher computational complexity for CU recursive detection and coding mode selection. However, the CU partitioning and the rapid selection of coding modes have been extensively researched. As a result, the computational load of recursive detection and coding mode selection has been significantly reduced[33, 34, 35, 36].
@@
In CMOS circuits, the overall circuit power consumption is primarily composed of three components: static power consumption, dynamic power consumption, and short-term power consumption, which can be represented as
@@
where $k$ is a constant in the dynamic power scaling model, determined by the supply voltage and the effective switching capacitance of the circuit[37].
@@ II-C2 Transmission Delay and Power Models
Transmission delay ${t_{trans}}$ refers to the total time it takes for a data packet to travel from the sender of the channel, through the transmission process, and ultimately reach the receiver. It mainly includes sending delay ${t_{send}}$, propagation delay ${t_{propa}}$, and buffer processing delay ${t_{buffer}}$, as illustrated in Fig. 2.
Figure 2: End-to-end delay composition of video coding and transmission.
@@ IV-B5 Comparative Analysis with HM RC Algorithm
IV-B6 Comparative Analysis of Subjective Performance
In this subsection, we also design a subjective experiment to validate the effectiveness of LyaRI in practical video coding and transmission applications. The experiment is based on the Coastguard sequence. Both LyaRI and the comparative method HM RC adopt the same initial coding parameters. Experimental results indicate that there is no visual artifacts or stalling in video decoding at the receiver using LyaRI. This is because LyaRI can adjust coding parameters to adapt to dynamic changes in the AtG channel. In contrast, HM RC method experiences visual artifacts at the $30$-th frame due to packet loss. Further, the packet loss results in a gradual degradation quality of subsequent video frames due to error propagation in video coding. For instance, Fig. 9(i) depicts the subjective quality of the $48$-th frame. It stems from the lack of a mechanism that adapts the source coding bitrate to the dynamic channel capacity, thus leading to packet loss. Besides, from Fig. 9, it can be observed that the subjective quality of videos encoded by both LyaRI and HM RC has declined compared to the original YUV video. The inefficiency of the HM coding process affects the video quality.
@@ IV-B6 Comparative Analysis of Subjective Performance
Figure 9: Subjective results for the Coastguard sequence.

### v5

@@ Abstract
This paper is concerned with unmanned aerial vehicle (UAV) video coding and transmission in scenarios such as aerial search and monitoring. Unlike existing methods of modeling UAV video source coding and channel transmission separately, we investigate the joint source-channel optimization issue for video coding and transmission. Particularly, we design eight-dimensional delay-power-rate-distortion models in terms of source coding and channel transmission and characterize the correlation between video coding and transmission, with which a joint source-channel optimization problem is formulated. Its objective is to minimize end-to-end distortion and UAV power consumption by optimizing fine-grained parameters related to UAV video coding and transmission. This problem is confirmed to be a challenging sequential-decision and non-convex optimization problem. We therefore decompose it into a family of repeated optimization problems by Lyapunov optimization and design an approximate convex optimization scheme with provable performance guarantees to tackle these problems. Based on the theoretical transformation, we propose a Lyapunov repeated iteration (LyaRI) algorithm. Both objective and subjective experiments are conducted to comprehensively evaluate the performance of LyaRI. The results indicate that, compared to its counterparts, LyaRI achieves better video quality and stability performance, with a 47.74% reduction in the variance of the obtained encoding bitrate.
@@ I Introduction
Benefiting from their distinctive advantages of low cost and high mobility, unmanned aerial vehicles (UAVs) have been increasingly adopted in more and more practical scenarios, such as geological disaster monitoring, forest fire detection, emergency communications, and emergency search and rescue [1, 2, 3]. In the application of UAVs, video acquisition and transmission constitute one of their pivotal missions. UAVs are capable of capturing targets through on-board cameras and subsequently transmitting coded and compressed videos via on-board data transmission modules. Rapid acquisition of dynamic information on site through UAV videos can save valuable time for emergency decision making. As a result, UAV video coding and transmission technology has emerged as a focal point of interest within both academia and industry [4, 5].
However, the timely and efficient transmission of UAV video streams is fraught with considerable challenges [6]. Firstly, a UAV network is bandwidth-constrained while the video is characterized by its substantial information and stringent timeliness. It is essential to mitigate information redundancy through video coding for UAV video transmission. Nevertheless, the computational complexity and energy expenditure associated with video coding are exceedingly high, presenting a significant challenge for UAVs that are subject to severe energy constraints. Secondly, the occurrence of congestion and packet loss, potentially leading to visual artifacts or stalling in video decoding at the receiver, may arise if the video coding bitrate surpasses the UAV channel capacity [7]. Further, UAV channel capacity is subject to dynamic fluctuations, and ensuring the match of UAV video coding bitrate with UAV channel capacity is a formidable challenge.
@@ I-A State of the Arts
All these studies assumed unlimited encoding time and power consumption. They also focused on the selection of video coding quantization parameters based on the R-D relationship and aimed to achieve the best encoding performance. However, the time and power consumption of video coding have been confirmed to significantly impact the video bitrate and distortion. In [14], the authors explored semi-extreme sparse coding of transformed residuals at medium and low bitrates to enhance video compression efficiency. The work in [15] indicated that video bitrate and distortion were primarily determined by the distribution characteristics of the transformed residuals. Transformed residuals were mainly influenced by motion estimation (ME) and quantization parameters. The accuracy of ME was closely related to the sum of absolute differences (SAD) of macroblocks. The number of calculations of macroblock SAD was proportional to the encoding complexity. Considering that the time and power consumption of video coding were increasing functions of encoding complexity, the authors in [15] deduced a quantitative relationship between video coding delay, power, rate, and distortion, and extended the traditional R-D model to a delay-power-rate-distortion (d-P-R-D) model. Based on this model, a new rate control (RC) method was proposed, aiming to minimize encoding distortion under constraints of rate, delay, and power. However, there are two aspects for improvement in this research. Firstly, the work is conducted for the earlier video coding standard advanced video coding (AVC). High efficiency video coding (HEVC) and versatile video coding (VVC), as the subsequent video coding standards[16], offer lower coding bitrates under the same video quality conditions. Therefore, it is necessary to further investigate the d-P-R-D model for more advanced standards. Secondly, although the end-to-end delay composition of video transmission systems was analyzed in [15], the proposed d-P-R-D model was only for video source coding. To achieve better end-to-end video transmission QoE, it is essential to design a joint source-channel video coding and transmission scheme.
@@
From the perspective of joint source-channel design, efficient video source coding can reduce the video bitrate while maintaining the same video quality. The reduction in video bitrate leads to decreasing transmit power consumption and transmission delay. However, efficient video source coding implies higher computational complexity, which result in increased power consumption and delay in video coding. Given the constraints of end-to-end latency, increasing encoding latency to achieve better compression performance may reduce the transmission latency budget, subsequently increasing the risk of distortion during transmission. Similarly, changes in encoding power consumption can affect the power consumption of video transmission under a given power constraint of video transmission system. Researchers have extensively studied joint source-channel optimization [17, 18, 19, 20]. For instance, the work in [17] proposed a new scheme for dynamically adjusting the bitrate based on channel signal-to-noise ratio (SNR) and image content. In [18], the authors presented a joint source-channel adaptive RC scheme for wireless image transmission based on image feature mapping, entropy awareness, and channel conditions. The authors in [19] introduced a novel video codec structure based on compressed sensing. They also investigated the minimization of system power consumption by jointly optimizing video coding bitrate and multi-path transmission bitrate under distortion and energy consumption constraints. The work in [20] investigated the trade-off between latency, power, and distortion and achieved deep cross-layer optimization of video coding, queuing control, and wireless transmission through a joint lossy compression and power allocation scheme. However, the aforementioned research focused on the field of wireless video transmission and did not specifically target UAV video transmission. To address this, the authors in [7] conducted cross-layer design research on UAV-based video streaming transmission. They proposed a quality of service (QoS) strategy that dynamically adjusted the transmission mode according to network load, latency, and packet loss rate. Nevertheless, they focused on the overall system design of multi-link parallel transmission, without delving into the issue of UAV video coding. Considering the dynamic nature and power constraints of UAVs, it is necessary to conduct joint resource allocation and control for UAV video coding and transmission under the constraints of end-to-end latency and power consumption. In this way, the QoE of video transmission at the receiving end can be improved.
@@ II-A AtG Channel Model
This paper primarily considers the application scenarios where UAVs are deployed for agricultural monitoring, forest fire detection, and emergency search and rescue. Particularly, the scenario investigated in this paper encompasses a UAV and an emergency communication vehicle (ECV). Given the potential for road traffic disruptions in areas such as mountainous and post-disaster zones, the ECV is deployed in proximity to the affected region to launch the UAV. The UAV is tasked with conducting surveillance and reconnaissance missions, capturing video footage for transmission, as illustrated in Fig. 1. The captured video is encoded and compressed by the UAV and then promptly delivered to the ECV. To facilitate the construction of a mathematical model for the video transmission process, the temporal domain is discretized into a sequence of time slots, denoted by $t=\{1,2,\ldots\}$. Considering some practical application requirements, the UAV is deployed to fly along a circular trajectory to extend the coverage range [21, 22, 23]. The two-dimensional (2-D) horizontal coordinate of the UAV at time slot $t$ is represented as $\bm{q_{uav}}\left(t\right)={[{x_{uav}}\left(t\right),{y_{uav}}\left(t\right)]^{\rm{T}}}$, and the position of the ECV is denoted as $\bm{q_{ecv}}={[{x_{ecv}},{y_{ecv}}]^{\rm{T}}}$. Consider that the UAV flies at a constant altitude $H$, and the altitude of the ECV is negligible compared to $H$. The horizontal projection distance between the UAV and the ECV at time slot $t$ can be expressed as $dis(t)=\sqrt{{{({x_{uav}}\left(t\right)-{x_{ecv}})}^{2}}+{{({y_{uav}}\left(t\right)-{y_{ecv}})}^{2}}}$.
@@ II-B1 Video Bitrate and Distortion Models
The research presented in this paper is based on an advanced standard, particularly HEVC. Although HEVC continues to employ the traditional block-based hybrid coding framework, which includes key modules such as prediction, transformation, quantization, entropy coding, and loop filtering, it introduces innovative coding techniques in each of these modules. These techniques include quadtree-based block partitioning, up to $35$ intra-prediction modes, inter-frame variable-sized discrete cosine transform (DCT), context-based adaptive binary arithmetic coding (CABAC), and sample adaptive offset (SAO) for loop filter compensation. Compared to AVC, HEVC significantly improves compression efficiency, achieving an approximately $50\%$ reduction in bitrate while maintaining the same video quality, making it highly suitable for video transmission in bandwidth-constrained networks. The VVC reference software VTM provides bit-rate savings of 44.4% on average when compared to HEVC; however, with more than 10.2 times higher computational costs on average [25, 26]. Considering the limited computational ability and power constraints of UAV, this paper constructs a d-P-R-D model for video coding based on HEVC.
@@
Besides, this paper adopts the IPPPP coding mode without loss of generality[27]. In the IPPPP coding mode, both encoding bitrate and distortion of P-frames in block-based hybrid video coding can be approximately represented as functions of the standard deviation of transformed residuals and the quantization step $Q$ [15]. Based on the research findings in literature[28], the functional expression $\sigma$ can be obtained through nonlinear regression analysis. This paper selects standard video test sequences CITY (CIF, 352*288) and Coastguard (CIF, 352*288) and employs the HM-16.20 encoder to obtain encoding data samples. In the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, since quantization parameter $QP$, search range $\lambda$, and reference frames $\chi$ are independent parameters, their individual impacts on $\sigma$ can be assessed separately. It should be noted that there is a one-to-one mapping relationship between the quantization parameter $QP$ mentioned in HM-16.20 and the quantization step $Q$ involved in the model. This paper first fixes $QP$ and $\chi$, and changes the value of $\lambda$ to obtain transformed residual data and calculate $\sigma$. Through nonlinear regression analysis, it has been found that $\sigma$ is approximately exponentially related to $\lambda$. Secondly, by fixing $\lambda$ and $\chi$, and changing $QP$, it can be concluded that $\sigma$ is approximately linearly related to $QP$. Additionally, it is observed that the rate of change of $\sigma$ versus $\chi$ is much smaller than that of $\sigma$ versus $QP$, as well as $\sigma$ versus $\lambda$. It indicates that $\chi$ has a relatively smaller impact on $\sigma$ compared to $QP$ and $\lambda$. To reduce the computational complexity, this paper assigns a given value ${\chi_{0}}=1$ to $\chi$[15]. Therefore, this paper adopts the following expression to calculate $\sigma$
@@ II-B2 Coding Delay and Power Models
In the research field of video coding, encoding complexity is an important indicator of quantization encoding delay. Compared to AVC, HEVC employs flexible coding units (CU) partitioning based on a quadtree structure and a richer set of coding modes. It results in relatively higher computational complexity for CU recursive detection and coding mode selection. However, the CU partitioning and the rapid selection of coding modes have been extensively researched. As a result, the computational load of recursive detection and coding mode selection has been significantly reduced [34, 35].
@@
In CMOS circuits, the overall circuit power consumption is primarily composed of three components: static power consumption, dynamic power consumption, and short-term power consumption, which can be represented as [15, 36]
@@
where $k$ is a constant in the dynamic power scaling model, determined by the supply voltage and the effective switching capacitance of the circuit.
@@ II-C2 Transmission Delay and Power Models
Transmission delay ${t_{trans}}$ refers to the total time it takes for a data packet to travel from the sender of the channel, through the transmission process, and ultimately reach the receiver. It mainly includes sending delay ${t_{send}}$, propagation delay ${t_{propa}}$, and buffer processing delay ${t_{buffer}}$.
@@ IV-B5 Comparative Analysis with HM RC Algorithm
(j) Original: Frame index = 121
(k) Original: Frame index = 170
(l) Original: Frame index = 200
(m) LyaRI: Frame index = 121
(n) LyaRI: Frame index = 170
(o) LyaRI: Frame index = 200
(p) HM RC: Frame index = 121
(q) HM RC: Frame index = 170
(r) HM RC: Frame index = 200
Figure 8: Subjective results for the Coastguard sequence and the City sequence.
IV-C Comparative Analysis of Subjective Performance
In this subsection, we also design a subjective experiment to validate the effectiveness of LyaRI in practical video coding and transmission applications. The experiment is based on the Coastguard sequence and the City sequence. Both LyaRI and the comparative method HM RC adopt the same initial coding parameters. Experimental results indicate that there is no visual artifacts or stalling in video decoding at the receiver using LyaRI. This is because LyaRI can adjust coding parameters to adapt to dynamic changes in the AtG channel. In contrast, HM RC method experiences visual artifacts at both the $30$-th frame of Coastguard and the $170$-th frame of City due to packet loss. Further, the packet loss results in a gradual degradation quality of subsequent video frames due to error propagation in video coding. For instance, Fig. 8(i) depicts the subjective quality of the $48$-th frame of Coastguard, and Fig. 8(r) the $200$-th frame of City. It stems from the lack of a mechanism that adapts the source coding bitrate to the dynamic channel capacity, thus leading to packet loss. Besides, from Fig. 8, it can be observed that the subjective quality of videos encoded by both LyaRI and HM RC has declined compared to the original YUV video. The inefficiency of the HM coding process affects the video quality.